OpenStudy (anonymous):

integrate e^(x^1/2) / x^(1/2) borns are 4 to 1

OpenStudy (anonymous):

Let u = x^1/2. Then du = (1/2)x^(-1/2) dx. So 2du = x^(-1/2) dx Then you’d be integrating 2*e^u du from u = 1 to u = 2.